OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 1, y = t2 - 1 a.y = x + 2, x ≥ 1 b.y = x - 2, x ≥ 1 c.y = x2 - 2, x ≥ 1 d. y = x2 + 2, x ≥ 1

OpenStudy (anonymous):

\[\begin{cases}x=t^2+1\\y=t^2-1\end{cases}\] Since both \(x\) and \(y\) contain a \(t^2\) term, it's easy to solve for \(t^2\): \[\begin{cases}t^2=x-1\\t^2=y+1\end{cases}\] So, you have \[x-1=y+1\\ ~~~~~~~~~~\vdots\]

OpenStudy (anonymous):

How would I simplify that into a standard form? @sinthsandgiggles

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Solve for \(y\) in terms of \(x\). Those are the provided answer choices, after all.

OpenStudy (anonymous):

Do u knw differentintion......?

OpenStudy (anonymous):

not at all @ashwinjohn3

OpenStudy (anonymous):

Uhm, no differentiation required. @grabyouroar, solve for \(y\) in the last equation I posted. You get \[y=x-2\]

OpenStudy (anonymous):

OHH okay now I understand. I was mistaken for something else. Thanks!@SithsAndGiggles

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Would you mind helping me with one more question? @SithsAndGiggles

OpenStudy (anonymous):

Ask away

OpenStudy (anonymous):

Eliminate the parameter. x = 7t, y = t + 4 Options are: y = 7x + 4 y = 7x - 4 y =( x/7) + 4 y = (x/7) - 4 @SithsAndGiggles

OpenStudy (anonymous):

This time, solving for \(t\) makes elimination simpler: \[x=7t~\Rightarrow~t=\frac{x}{7}\\ y=t+4~\Rightarrow~t=y-4\] So you're left with \[\frac{x}{7}=y-4\]

OpenStudy (anonymous):

So my end product will be a fraction? @SithsAndGiggles

OpenStudy (anonymous):

WAIT nvm lol

OpenStudy (anonymous):

Solving for \(y\), you're left with \[y=\frac{x}{7}+4\] (the third answer choice) The answer itself isn't really a fraction (more of a mixed number).

OpenStudy (anonymous):

okay *face palm how do you know all this!?

OpenStudy (anonymous):

Paying attention in class? Idk, this is pretty basic algebra haha

OpenStudy (anonymous):

math's not my thing man! haha thanks again @SithsAndGiggles